Ok, I'm not good at word problems and I need a little help if possible.
A ladder is leaning against a building. The distance from the bottom of the ladder to the building is 8 ft less than the length of the ladder. How high up the side of the building is the top of the ladder if that distance is 4 ft less than the length of the ladder.
The ladder reached ___ feet up the wall.

Answer by checkley75(3666)

(X-8)^2+(X-4)^2=X^2
X^2-16X+64+X^2-8X+16=X^2
2X^2-24X+80-X^2=0
X^2-24X+80=0
(X-4)(X-20)=0
X-20=0
X=20 ANS. FOR THE LENGTH OF THE LADDER.
PROOF:
(20-8)^2+(20-4)^2=20^2
12^2+16^2=400
144+1256=400
400=400